# Simultaneous Motion And Noise Estimation with Event Cameras

### v1

@@ Abstract
Event cameras are emerging vision sensors, whose noise is challenging to characterize. Existing denoising methods for event cameras consider other tasks such as motion estimation separately (i.e., sequentially after denoising). However, motion is an intrinsic part of event data, since scene edges cannot be sensed without motion. This work proposes, to the best of our knowledge, the first method that simultaneously estimates motion in its various forms (e.g., ego-motion, optical flow) and noise. The method is flexible, as it allows replacing the 1-step motion estimation of the widely-used Contrast Maximization framework with any other motion estimator, such as deep neural networks. The experiments show that the proposed method achieves state-of-the-art results on the E-MLB denoising benchmark and competitive results on the DND21 benchmark, while showing its efficacy on motion estimation and intensity reconstruction tasks. We believe that the proposed approach contributes to strengthening the theory of event-data denoising, as well as impacting practical denoising use-cases, as we release the code upon acceptance. Project page: https://github.com/tub-rip/ESMD
@@ 1 Introduction
Event cameras are emerging vision sensors that overcome challenges of conventional cameras (e.g., motion blur, limited dynamic range, data and power efficiency), yet they suffer from a considerable amount of noise due to their novelty and operation in low-power (transistor subthreshold) conditions [19]. Since event cameras are suitable for many computer vision tasks, especially motion-related tasks, it is paramount to classify event data that is related to motion (i.e., signal) and that is not (i.e., noise). However, denoising event data is a challenging problem, since noise properties are not fully characterized to date, and it is not feasible to define ground-truth (GT) noise labels in real-world data recordings. Previous work either ($i$) leveraged simulation to generate pure signal (events from edge motion under constant illumination) and noise, or ($ii$) prepared real event data that was expected to be clean, by aggressively removing noise-like events with existing filters and using bright illumination that minimized noise occurrence.
@@ 2 Related Work
Denoising is a fundamental problem in event cameras since there are a lot of leakage events [24], and event-processing algorithms may suffer in scenarios with high noise [3]. Removing event noise has a double effect: it reduces the load of subsequent processing stages (as the load is proportional to the number of events) and improves their output (as the input data becomes of higher quality). In this work, we focus on noise due to the leakage inside the camera hardware, such as background activity (BA) events [35]. Events due to flickering or active lights can be treated as “noise” for methods that consider passive sensing, which are a large majority in computer vision, and are therefore not our focus.
@@
Our work builds upon a series of successful methods on motion estimation and their applications. Event-based motion estimation is a paramount topic since event cameras naturally respond to motion in the scenes [19]. In particular, Contrast Maximization (CMax) [17, 49] is a state-of-the-art framework for motion estimation, which has various applications, such as rotational motion [15, 23, 44, 41, 31], optical flow [62, 43, 47, 27], intensity reconstruction [58], and SLAM [25, 22, 54]. We investigate the connection and potential benefits of applying the CMax framework to denoising during the estimation process (i.e., simultaneously), demonstrating motion estimation improvements as well as competitive denoising accuracy. A prior work, “ProgressiveMotionSeg” [8], has extended the CMax-based segmentation method [50] with extra weights that accounted for how much each event contributed to estimated motion clusters. However, [8] mainly focuses on the task of motion segmentation using low-DOF warps (e.g., 2-DOF feature flow) and does not discuss denoising efficacy or evaluate on standard denoising benchmarks. In contrast, our method supports various types of motion estimation (both low-DOF and high-DOF warps like optical flow), and considers that noise shall not be used to estimate motion, as from first principles we deem noise and motion to be uncorrelated. We propose signal-noise classifiers and show ranking invariance, which results in state-of-the-art denoising accuracy and improvements on motion estimation from noisy data.
@@ 3 Methodology
The following sections present the CMax framework (Sec. 3.1), the proposed signal/noise classifier (Sec. 3.2), and the simultaneous denoiser and motion estimator (Sec. 3.3).
@@ 3.1 Contrast Maximization
$\vspace{-0.5ex}e_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})\;\,\mapsto\;\,e^{\prime}_{k}\doteq(\mathbf{x}^{\prime}_{k},t_{\text{ref}},p_{k}).\vspace{-0.5ex}$ (1)
@@
$\vspace{-0.5ex}\textstyle I(\mathbf{x};\mathcal{E}^{\prime}_{t_{\text{ref}}},\boldsymbol{\theta})\doteq\sum_{k=1}^{N_{e}}\delta(\mathbf{x}-\mathbf{x}^{\prime}_{k}),\vspace{-0.5ex}$ (2)
@@
GasTank (E-MLB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 2: Qualitative result of denoising and flow estimation on E-MLB and DND21 data. The driving sequence has $5$-Hz artificial noise. (b) is an alternative visualization of (a) with the same grayscale coding as (d) for ease of comparison. Flow color is given in Fig. 1.
@@ 3.2 Denoising Based on Local Contrast
As Fig. 1 shows, our method is iterative. In each iteration, motion estimation is refined and events are classified into signal $\mathcal{E}_{s}$ or noise $\mathcal{E}_{n}$. Without labels or additional information, events rely on each other to realize this classification. To this end, we compute a score $c_{k}$ for each event, rank the events according to their score, and threshold at some value $T$. The threshold is computed by assuming the amount or ratio of noisy events is known $\eta$, i.e., $T\equiv T(\eta)$, which allows us to have high-level control over the denoising process. Hence we classify events into signal (the $\tau\doteq(1-\eta)$ percentage of events with highest scores) and noise:
@@
We studied different choices for the per-event score $c_{k}$ based on the agreement among warped events $\mathcal{E}^{\prime}=\mathcal{E}^{\prime}_{s}\cup\mathcal{E}^{\prime}_{n}$, through their IWEs, and concluded on two. Letting $I^{s}_{k}\equiv I^{s}_{k}(\mathbf{x}^{\prime}_{k}),I^{n}_{k}\equiv I^{n}_{k}(\mathbf{x}^{\prime}_{k})$ and $I_{k}\equiv I(\mathbf{x}^{\prime}_{k})=I^{s}_{k}+I^{n}_{k}$ be the IWE values of previously classified signal events, noise events and all events, respectively, the two choices are: ($i$) ${}^{1}c_{k}\doteq I_{k}$, and ($ii$) the ratio of signal events ${}^{2}c_{k}\doteq I^{s}_{k}/I_{k}$ (probabilistic viewpoint). In fact, both are equivalent: rule ${}^{1}c_{k}>T_{1}$ is the same as rule ${}^{2}c_{k}>T_{2}$ if $T_{2}=1-I^{n}_{k}/T_{1}$ and the event noise $I^{n}_{k}$ is known. We use ${}^{1}c_{k}$, the number of warped events (also called local contrast [18, 50]), because it is the simplest to interpret.
@@ 3.3 Alternating Optimization
Classifying events into signal and noise requires knowledge of the true motion, and estimation of the true motion (e.g., using CMax) requires knowledge of the signal events because, by definition, noise events carry no information about motion. Hence, it is a chicken-and-egg problem.
We approach its solution iteratively, as shown in Fig. 1, by combining motion estimation (Sec. 3.1) and S/N classification (Sec. 3.2). Assuming some initialization, the current signal events $\mathcal{E}_{\text{signal}}$ are used to estimate the motion via CMax (1 step is enough); then all events are warped to compute the IWE and the per-event scores $c_{k}$, from which the sets (3) are recomputed, in preparation for the next iteration.
In practice, we initialize with a random split of $\mathcal{E}$ into $\mathcal{E}_{\text{signal}}$ and $\mathcal{E}_{\text{noise}}$, update $\mathcal{E}_{\text{signal}}$ at every warp during the optimization iteration until the process converges (as flagged by the convergence of the motion parameters in CMax).
@@ 4 Experiments
E-MLB (Day) E-MLB (Night) DND21 ND1 ND4 ND16 ND64 ND1 ND4 ND16 ND64 Model-based Raw $0.821$ $0.824$ $0.815$ $0.786$ $0.89$ $0.824$ $0.786$ $0.768$ $0.869$ BAF [9] $0.861$ $0.869$ $0.876$ $0.89$ $0.946$ $0.973$ $0.992$ $0.942$ $0.92$ KNoise [30] $0.846$ $0.837$ $0.83$ $0.807$ $0.954$ $0.956$ $0.871$ $0.817$ $0.887$ DWF [24] $0.878$ $0.876$ $0.866$ $0.865$ $0.923$ $0.962$ $0.988$ $0.932$ $0.905$ EvFlow [53] $0.848$ $0.878$ $0.868$ $0.833$ $0.969$ $0.983$ $0.889$ $0.797$ $1.006$ Ynoise [14] $0.866$ $0.863$ $0.857$ $0.821$ $1.009$ $0.943$ $0.875$ $0.792$ $0.966$ TS [33] $0.877$ $0.887$ $0.87$ $0.837$ $1.033$ $0.944$ $0.886$ $0.797$ $0.985$ IETS [5] $0.772$ $0.785$ $0.777$ $0.753$ $0.950$ $0.823$ $0.804$ $0.711$ $0.900$ GEF [11] $1.051$ $0.938$ $0.935$ $0.927$ $1.027$ $0.955$ $0.946$ $0.935$ $0.932$ Ours 0.938 254 384 292 974 70.93825438429297470.938\,254\,384\,292\,974\,70.938 254 384 292 974 7 0.958 464 335 277 024 80.95846433527702480.958\,464\,335\,277\,024\,80.958 464 335 277 024 8 0.985 896 450 122 991 80.98589645012299180.985\,896\,450\,122\,991\,80.985 896 450 122 991 8 0.950 081 880 240 459 60.95008188024045960.950\,081\,880\,240\,459\,60.950 081 880 240 459 6 1.036 831 015 641 0911.0368310156410911.036\,831\,015\,641\,0911.036 831 015 641 091 0.961 426 679 598 1070.9614266795981070.961\,426\,679\,598\,1070.961 426 679 598 107 0.944 527 953 022 538 40.94452795302253840.944\,527\,953\,022\,538\,40.944 527 953 022 538 4 0.932 362 037 844 1260.9323620378441260.932\,362\,037\,844\,1260.932 362 037 844 126 0.992 069 121 10.99206912110.992\,069\,121\,10.992 069 121 1 Learning EventZoom [12] $0.996$ $0.988$ $0.996$ $0.97$ $1.055$ $1.007$ $1.01$ $0.988$ $1.059$ EDnCNN [4] $0.887$ $0.908$ $0.903$ $0.912$ $1.001$ $1.024$ $1.079$ $1.086$ $0.977$ MLPF [24] $0.851$ $0.855$ $0.846$ $0.84$ $0.926$ $0.928$ $0.91$ $0.906$ $0.944$ EDformer [28] $0.952$ $0.955$ $0.956$ $0.942$ $1.048$ $1.019$ $1.076$ $1.099$ $1.041$
@@
1Hz 5Hz 10Hz hotel driving hotel driving hotel driving Model-based KNoise [30] $0.6773$ $0.6296$ $0.6703$ $0.6235$ $0.6413$ $0.6142$ DWF [24] $0.9268$ $0.7409$ $0.8620$ $0.6901$ $0.7958$ $0.6563$ BAF [9] $0.9535$ $0.8479$ $0.8916$ $0.7930$ $0.8366$ $0.7479$ Ynoise [14] $0.9690$ $0.9409$ $0.9234$ $0.9093$ $0.8987$ $0.8800$ TS [33] $0.9716$ $0.9307$ $0.9606$ $0.9270$ $0.9620$ $0.9202$ Ours 1.013 802 194 317 176 21.01380219431717621.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.88211969313369480.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.962 515 815 619 444 40.96251581561944440.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.85539848463811620.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.960 564 724 999 644 20.96056472499964420.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.8362735551022360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] $0.9573$ $0.8873$ $0.9365$ $0.8748$ $0.9006$ $0.874$ MLPF [24] $0.9704$ $0.8887$ $0.9704$ $0.8845$ $0.9634$ $0.8761$ EDformer [28] $0.9928$ $0.9541$ $0.9845$ $0.9424$ $0.9699$ $0.9264$
@@
Figure 4: Comparison of various denoising methods on the two DND21 sequences (driving and hotel) at two different noise levels.
4.1 Datasets, Metrics, and Hyper-parameters
Datasets. E-MLB [10] is a large-scale, de-facto dataset to benchmark denoising. It consists of $100$ sequences with 4 different brightness levels using neutral-density (ND) filters on a DAVIS346 camera ($346\times 260$ px) during day and night.
Historically, the two sequences (hotel and driving) of the DND21 dataset [24] have been widely used for denoising evaluation [24, 10, 28]. Both sequences are recorded with a DAVIS346 camera [52]. They are regarded as “signal” data, as the sequences have been aggressively filtered offline. Combined with other pure BA noise sequences, the dataset has event-wise noise annotations that are useful for evaluation. We use the above sequences with different noise rates, $1,3,5,7,10$ Hz per pixel, following prior work [28].
The ECD dataset [39] is a standard dataset for various tasks including camera ego-motion estimation [15, 60, 46, 40, 59, 48]. Using a DAVIS240C camera ($240\times 180$ px [7]), each sequence provides events, frames, calibration information, IMU data, and ground truth (GT) camera poses (at $200$ Hz). We use boxes_rotation and dynamic_rotation sequences to evaluate rotational motion estimation accuracy.
@@ 4.1 Datasets, Metrics, and Hyper-parameters
Metrics. The metrics used for denoising assessment are: receiver operating characteristic (ROC) curve, area under the curve (AUC) and Mean Event Structural Ratio (MESR). Given signal and noise GT labels, the ROC shows the True-Positive Rate and False-Positive Rate over several S/N ratios. The AUC of the ROC becomes closer to $1$, the better the denoiser works. The MESR is a denoising metric that does not require GT labels [10]. We follow previous work to fix the number of signal events to 30000 [28, 10], as well as $M=20000$ [10, Eq.10], to discount the difference in the number of signal events from various denoising methods. The higher, the better.
We use the RMS error for angular velocity estimation and the EPE for optical flow estimation, following previous works [41, 23, 25, 49], as well as the FWL metric [51], which is the relative variance (sharpness) of the IWE with respect to that of the identity warp.

### v2

@@ Abstract
Event cameras are emerging vision sensors whose noise is challenging to characterize. Existing denoising methods for event cameras are often designed in isolation and thus consider other tasks, such as motion estimation, separately (i.e., sequentially after denoising). However, motion is an intrinsic part of event data, since scene edges cannot be sensed without motion. We propose, to the best of our knowledge, the first method that simultaneously estimates motion in its various forms (e.g., ego-motion, optical flow) and noise. The method is flexible, as it allows replacing the one-step motion estimation of the widely-used Contrast Maximization framework with any other motion estimator, such as deep neural networks. The experiments show that the proposed method achieves state-of-the-art results on the E-MLB denoising benchmark and competitive results on the DND21 benchmark, while demonstrating effectiveness across motion estimation and intensity reconstruction tasks. Our approach advances event-data denoising theory and expands practical denoising use-cases via open-source code. Project page: https://github.com/tub-rip/ESMD
{textblock}
11(2.5, 0.6)
This paper has been accepted for publication at the
IEEE International Conference on Computer Vision (ICCV), Honolulu, 2025. ©IEEE
@@ 1 Introduction
Event cameras are emerging vision sensors that overcome shortcomings of conventional cameras (e.g., motion blur, limited dynamic range, data and power efficiency), yet they suffer from a considerable amount of noise due to their novelty and operation in low-power (transistor subthreshold) conditions [19, 36]. Since event cameras are suitable for many computer vision tasks, especially motion-related tasks, it is paramount to classify event data that is related to motion (i.e., signal) and that is not (i.e., noise). However, denoising event data is a challenging problem because noise properties have yet to be fully characterized [24, 22] and it is infeasible to define ground-truth (GT) noise labels in real-world data recordings. Previous works either leveraged simulation or prepared real event data and aggressively suppressed noise occurrence.
@@ 2 Related Work
Denoising is a fundamental problem in event cameras since there are a lot of leakage events [24], and event-processing algorithms may suffer in scenarios with high noise [3]. Removing event noise has a double effect: it reduces the load of subsequent processing stages (as the load is proportional to the number of events) and improves their output (as the input data becomes of higher quality). In this work, we focus on noise due to the leakage inside the camera hardware, such as background activity (BA) events [36]. Events due to flickering or active lights are therefore not our priority.
@@
Our work builds upon a series of successful methods for motion estimation and their applications. Event-based motion estimation is an important topic since event cameras naturally respond to motion in the scenes [19]. In particular, Contrast Maximization (CMax) [17, 50] is a state-of-the-art framework for motion estimation, which has various applications, such as rotational motion estimation [15, 23, 45, 42, 32], optical flow estimation [64, 44, 48, 28, 27], and motion segmentation [51, 60, 57]. We investigate the connection and potential benefits of applying the CMax framework to denoising during the estimation process (i.e., simultaneously), demonstrating motion estimation improvements as well as competitive denoising accuracy.
A prior work, “ProgressiveMotionSeg” [8], has extended the CMax-based segmentation method [51] with extra weights that accounted for how much each event contributed to the estimated motion clusters. However, [8] mainly focuses on the task of motion segmentation using low degrees-of-freedom (DOF) warps (e.g., 2-DOF feature flow) and does not discuss denoising efficacy or evaluate on standard denoising benchmarks. In contrast, our method supports various types of motion estimation (both low-DOF and high-DOF warps, such as optical flow), and considers that noise shall not be used to estimate motion, as from first principles we deem noise and motion to be uncorrelated.
@@ 3.1 Contrast Maximization
$\vskip-2.58334pte_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})\;\,\mapsto\;\,e^{\prime}_{k}\doteq(\mathbf{x}^{\prime}_{k},t_{\text{ref}},p_{k}).\vskip-0.86108pt$ (1)
@@
$\vskip-2.58334pt\textstyle I(\mathbf{x};\mathcal{E}^{\prime}_{t_{\text{ref}}},\boldsymbol{\theta})\doteq\sum_{k=1}^{N_{e}}\delta(\mathbf{x}-\mathbf{x}^{\prime}_{k}),\vskip-0.86108pt$ (2)
@@
GasTank (E-MLB)
@@
Figure 2: Qualitative result of denoising and flow estimation on E-MLB and DND21. The driving sequence has 555-Hz artificial noise. (b) is an alternative visualization of (a) with the same grayscale coding as (d). Flow is visualized on all events and flow color is given in Fig. 1.
@@ 3.2 Denoising Based on Local Contrast
As Fig. 1 shows, our method is iterative. In each iteration, motion estimation is refined and events are classified into signal $\mathcal{E}_{s}$ or noise $\mathcal{E}_{n}$. Without labels or additional information, events rely on each other to realize this classification. To this end, we compute a score $c_{k}$ for each event, rank the events according to their scores, and threshold at some value $T$. The threshold is computed by assuming the amount or ratio of noisy events is known $\eta$, i.e., $T\equiv T(\eta)$, which allows us to have high-level control over the denoising process. Hence we classify events into signal (the $\tau\doteq(1-\eta)$ percentage of events with highest scores) and noise:
@@
We studied different choices for the per-event score $c_{k}$ based on the agreement among warped events $\mathcal{E}^{\prime}=\mathcal{E}^{\prime}_{s}\cup\mathcal{E}^{\prime}_{n}$, through their IWEs, and concluded on two, as follows. Letting $I^{s}_{k}\equiv I^{s}_{k}(\mathbf{x}^{\prime}_{k}),I^{n}_{k}\equiv I^{n}_{k}(\mathbf{x}^{\prime}_{k})$ and $I_{k}\equiv I(\mathbf{x}^{\prime}_{k})=I^{s}_{k}+I^{n}_{k}$ be the IWE values of previously classified signal events, noise events and all events, respectively, the two choices are: ($i$) ${}^{1}c_{k}\doteq I_{k}$, and ($ii$) the ratio of signal events ${}^{2}c_{k}\doteq I^{s}_{k}/I_{k}$ (probabilistic viewpoint). In fact, both are equivalent: rule ${}^{1}c_{k}>T_{1}$ is the same as rule ${}^{2}c_{k}>T_{2}$ if $T_{2}=1-I^{n}_{k}/T_{1}$ and the event noise $I^{n}_{k}$ is known. We use ${}^{1}c_{k}$, the number of warped events (also called local contrast [18, 51]), because it is the simplest to interpret.
@@ 3.3 Alternating Optimization
Classifying events into signal and noise requires knowledge of the true motion, and estimation of the true motion (e.g., using CMax) requires knowledge of the signal events because, by definition, noise events carry no information about motion. Hence, it is a circular dependency problem.
We approach its solution iteratively, as shown in Fig. 1, by combining motion estimation (Sec. 3.1) and S/N classification (Sec. 3.2). Assuming some initialization, the current signal events $\mathcal{E}_{\text{signal}}$ are used to estimate the motion via CMax (1 step is enough); then all events are warped to compute the IWE and the per-event scores $c_{k}$, from which the sets (3) are recomputed, in preparation for the next iteration. In practice, we initialize with a random split of $\mathcal{E}$ into $\mathcal{E}_{\text{signal}}$ and $\mathcal{E}_{\text{noise}}$, update $\mathcal{E}_{\text{signal}}$ at every warp during the optimization iteration until the process converges (as flagged by the convergence of the motion parameters in CMax).
@@ 4 Experiments
4.1 Datasets, Metrics, and Hyper-parameters
Datasets. E-MLB [10] is a large-scale, de-facto dataset to benchmark denoising. It consists of 100100100 sequences with 4 different brightness levels using neutral-density (ND) filters on a DAVIS346 camera (346×260346$\times$ 260346 × 260 px) during day and night.
Historically, the two sequences (hotel and driving) of the DND21 dataset [24] have been widely used for denoising evaluation [24, 10, 29]. Both sequences are recorded with a DAVIS346 camera [53]. They are regarded as “signal” data, as the sequences have been aggressively filtered offline. Combined with other pure BA noise sequences, the dataset has event-wise noise annotations that are useful for evaluation. We use the above sequences with different noise rates, 1,3,5,7,101,3,5,7,101 , 3 , 5 , 7 , 10 Hz per pixel, following prior work [29].
E-MLB (Day) E-MLB (Night) DND21 ND1 ND4 ND16 ND64 ND1 ND4 ND16 ND64 Model-based Raw 0.8210.8210.821 0.8240.8240.824 0.8150.8150.815 0.7860.7860.786 0.890.890.89 0.8240.8240.824 0.7860.7860.786 0.7680.7680.768 0.8690.8690.869 BAF [9] 0.8610.8610.861 0.8690.8690.869 0.8760.8760.876 0.890.890.89 0.9460.9460.946 0.9730.9730.973 0.9920.9920.992 0.9420.9420.942 0.920.920.92 TS [34] 0.8770.8770.877 0.8870.8870.887 0.870.870.87 0.8370.8370.837 1.0331.0331.033 0.9440.9440.944 0.8860.8860.886 0.7970.7970.797 0.9850.9850.985 KNoise [31] 0.8460.8460.846 0.8370.8370.837 0.830.830.83 0.8070.8070.807 0.9540.9540.954 0.9560.9560.956 0.8710.8710.871 0.8170.8170.817 0.8870.8870.887 EvFlow [54] 0.8480.8480.848 0.8780.8780.878 0.8680.8680.868 0.8330.8330.833 0.9690.9690.969 0.9830.9830.983 0.8890.8890.889 0.7970.7970.797 1.0061.0061.006 IETS [5] 0.7720.7720.772 0.7850.7850.785 0.7770.7770.777 0.7530.7530.753 0.9500.9500.950 0.8230.8230.823 0.8040.8040.804 0.7110.7110.711 0.9000.9000.900 Ynoise [14] 0.8660.8660.866 0.8630.8630.863 0.8570.8570.857 0.8210.8210.821 1.0091.0091.009 0.9430.9430.943 0.8750.8750.875 0.7920.7920.792 0.9660.9660.966 GEF [11] 1.0511.0511.051 0.9380.9380.938 0.9350.9350.935 0.9270.9270.927 1.0271.0271.027 0.9550.9550.955 0.9460.9460.946 0.9350.9350.935 0.9320.9320.932 DWF [24] 0.8780.8780.878 0.8760.8760.876 0.8660.8660.866 0.8650.8650.865 0.9230.9230.923 0.9620.9620.962 0.9880.9880.988 0.9320.9320.932 0.9050.9050.905 Ours 0.938 254 384 292 974 70.938\,254\,384\,292\,974\,70.938 254 384 292 974 7 0.958 464 335 277 024 80.958\,464\,335\,277\,024\,80.958 464 335 277 024 8 0.985 896 450 122 991 80.985\,896\,450\,122\,991\,80.985 896 450 122 991 8 0.950 081 880 240 459 60.950\,081\,880\,240\,459\,60.950 081 880 240 459 6 1.036 831 015 641 0911.036\,831\,015\,641\,0911.036 831 015 641 091 0.961 426 679 598 1070.961\,426\,679\,598\,1070.961 426 679 598 107 0.944 527 953 022 538 40.944\,527\,953\,022\,538\,40.944 527 953 022 538 4 0.932 362 037 844 1260.932\,362\,037\,844\,1260.932 362 037 844 126 0.992 069 121 10.992\,069\,121\,10.992 069 121 1 Learning EDnCNN [4] 0.8870.8870.887 0.9080.9080.908 0.9030.9030.903 0.9120.9120.912 1.0011.0011.001 1.0241.0241.024 1.0791.0791.079 1.0861.0861.086 0.9770.9770.977 EventZoom [12] 0.9960.9960.996 0.9880.9880.988 0.9960.9960.996 0.970.970.97 1.0551.0551.055 1.0071.0071.007 1.011.011.01 0.9880.9880.988 1.0591.0591.059 MLPF [24] 0.8510.8510.851 0.8550.8550.855 0.8460.8460.846 0.840.840.84 0.9260.9260.926 0.9280.9280.928 0.910.910.91 0.9060.9060.906 0.9440.9440.944 EDformer [29] 0.9520.9520.952 0.9550.9550.955 0.9560.9560.956 0.9420.9420.942 1.0481.0481.048 1.0191.0191.019 1.0761.0761.076 1.0991.0991.099 1.0411.0411.041
@@ 4.1 Datasets, Metrics, and Hyper-parameters
1Hz 5Hz 10Hz hotel driving hotel driving hotel driving Model-based BAF [9] 0.95350.95350.9535 0.84790.84790.8479 0.89160.89160.8916 0.79300.79300.7930 0.83660.83660.8366 0.74790.74790.7479 TS [34] 0.97160.97160.9716 0.93070.93070.9307 0.96060.96060.9606 0.92700.92700.9270 0.96200.96200.9620 0.92020.92020.9202 KNoise [31] 0.67730.67730.6773 0.62960.62960.6296 0.67030.67030.6703 0.62350.62350.6235 0.64130.64130.6413 0.61420.61420.6142 Ynoise [14] 0.96900.96900.9690 0.94090.94090.9409 0.92340.92340.9234 0.90930.90930.9093 0.89870.89870.8987 0.88000.88000.8800 DWF [24] 0.92680.92680.9268 0.74090.74090.7409 0.86200.86200.8620 0.69010.69010.6901 0.79580.79580.7958 0.65630.65630.6563 Ours 1.013 802 194 317 176 21.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.962 515 815 619 444 40.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.960 564 724 999 644 20.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] 0.95730.95730.9573 0.88730.88730.8873 0.93650.93650.9365 0.87480.87480.8748 0.90060.90060.9006 0.8740.8740.874 MLPF [24] 0.97040.97040.9704 0.88870.88870.8887 0.97040.97040.9704 0.88450.88450.8845 0.96340.96340.9634 0.87610.87610.8761 EDformer [29] 0.99280.99280.9928 0.95410.95410.9541 0.98450.98450.9845 0.94240.94240.9424 0.96990.96990.9699 0.92640.92640.9264
@@
Figure 4: Comparison of denoising methods on the two DND21 sequences at two different noise levels.
The ECD dataset [40] is a standard dataset for various tasks including camera ego-motion estimation [15, 62, 47, 41, 61, 49]. Using a DAVIS240C camera (240×180240$\times$ 180240 × 180 px [7]), each sequence provides events, frames, calibration information, IMU data, and ground truth (GT) camera poses (at 200200200 Hz). We use boxes_rotation and dynamic_rotation sequences to evaluate rotational motion estimation accuracy.
@@
Metrics. The metrics used for denoising assessment are: receiver operating characteristic (ROC) curve, area under the curve (AUC) and Mean Event Structural Ratio (MESR). Given signal and noise GT labels, the ROC shows the True-Positive Rate and False-Positive Rate over several S/N ratios. The AUC of the ROC becomes closer to 111, the better the denoiser works. The MESR is a denoising metric that does not require GT labels [10]. We follow previous work to fix the number of signal events to 30000 [29, 10], as well as $M=20000$ [10, Eq.10], to discount the difference in the number of signal events from various denoising methods. The higher, the better. Furthermore, we use the RMS error for angular velocity estimation and the EPE for optical flow estimation, following previous works [42, 23, 25, 50], as well as the FWL metric [52], which is the relative variance (sharpness) of the IWE with respect to that of the identity warp. Further details about the experimental settings (e.g., hyper-parameters) are in the supplementary.